A current of 3 A flows through a resistor for 6 minutes. Determine the power dissipated in the resistor if 3V existed across the resistor.

1
Expert's answer
2021-11-04T18:04:12-0400

Power

"P=i^2R"

"R=\frac{V}{I}=\frac{3}{3}=1A"

"P=3^2\times1=9W"

Work(W)="Pt=9\times 360=3240J"